A passanger train speed is 60mi/h and a freight train speed is 40 mi/h. The passenger train travel same diastance in 1.5 h less time than the freight train. How long does each train take to make the trip?
Answer by checkley77(12844)

D=RT
D=60(T-1.5)
D=40T
60(T-1.5)=40T
60T-90=40T
60T-40T=90
20T=90
T=90/20
T=4.5 HOURS FOR THE FREIGHT TRAIN.
4.5-1.5=3 hours for the passenger train.
Proof:
60*3=40*4.5
180=180